### v1

# Authentic Discrete Diffusion Model

Abstract
We propose an Authentic Discrete Diffusion (ADD) framework that fundamentally redefines prior pseudo-discrete approaches by preserving core diffusion characteristics directly in the one-hot space through a suite of coordinated mechanisms. Unlike conventional “pseudo” discrete diffusion (PDD) methods, ADD reformulates the diffusion input by directly using float-encoded one-hot class data, without relying on diffusing in the continuous latent spaces or masking policies. At its core, a timestep-conditioned cross-entropy loss is introduced between the diffusion model’s outputs and the original one-hot labels. This synergistic design establishes a bridge between discriminative and generative learning. Our experiments demonstrate that ADD not only achieves superior performance on classification tasks compared to the baseline, but also exhibits excellent text generation capabilities on Image captioning. Extensive ablations validate the measurable gains of each component.
[…]
11footnotetext: * Equal contribution.22footnotetext: † Corresponding author.
1 Introduction
Diffusion models have emerged as a powerful class of generative methods, achieving state-of-the-art performance in continuous domains by modeling data generation as a reverse denoising process (Ho et al., 2020). These models operate in continuous-valued vector spaces and are typically trained using mean squared error (MSE) to reconstruct data from Gaussian noise. However, this formulation is inherently incompatible with discrete signals, whose categorical structure and non-Euclidean geometry introduce significant challenges for both optimization stability and sampling fidelity.
Efforts to extend diffusion models to discrete domains have generally followed two unsatisfactory paths. One line of work maps discrete variables into continuous latent spaces via embedding, but these approaches often suffer from degraded generation quality and unstable training (Gong et al., 2023). Another line frames “discrete diffusion” as a masked modeling task, where noise is simulated through random token masking and the model is trained to reconstruct the original input (Google DeepMind, 2025; Nie et al., 2025; Wu et al., 2025; Ye et al., 2025). However, such approaches effectively replicate masked language modeling as popularized by BERT (Devlin et al., 2019), and do not satisfy the formal definition of a diffusion process (Sohl-Dickstein et al., 2015). Moreover, their empirical performance often remains inferior to strong autoregressive baselines. For this reason, we refer to such methods as “pseudo” discrete diffusion (PDD).
In this work, we introduce the Authentic Discrete Diffusion (ADD) model—a framework that preserves the defining properties of diffusion while operating directly in discrete spaces. ADD starts from Gaussian-corrupted one-hot vectors and iteratively denoises them over multiple steps following a well-defined noise schedule. At each step during inference, the model predicts a clean one-hot vector by first applying an $\arg\max$ to its output probabilities, then converting the result into a one-hot representation. This vector is fed back into the next iteration after adding noise with a reduced coefficient. This autoregressive-style feedback loop enables progressive refinement, allowing the model to eliminate uncertainty over time and converge toward semantically precise, categorical outputs.
Figure 1: Comparison between (a) traditional continuous-space diffusion, (b) masking-based pseudo-discrete diffusion (i.e., “pseudo” discrete diffusion (PDD)), and (c) our proposed Authentic Discrete Diffusion (ADD) framework. In (c), the left panel illustrates single-token discrete generation (e.g., classification), while the right panel illustrates multi-token discrete generation (e.g., text generation). Continuous diffusion models operate in Gaussian space, performing noise prediction and MSE-based reconstruction. Pseudo-discrete models mimic diffusion through masked token recovery. In contrast, ADD begins from Gaussian-corrupted one-hot vectors and performs true denoising entirely in the one-hot space. At each step, the model applies an $\arg\max$ to produce discrete predictions, converts them into one-hot vectors, and feeds them into the next iteration after adding noise with a reduced coefficient. This autoregressive-style refinement yields stable and efficient discrete-space diffusion, achieving accurate categorical predictions in only a few steps.
Crucially, ADD applies timestep-conditioned cross-entropy loss to directly enforce correspondence between the predicted and target one-hot vectors, while preventing the diffusion network from over-relying on conditioning at the expense of ignoring the input during training. This avoids the smoothing effects of MSE losses and ensures that outputs respect the mutually exclusive nature of discrete categories. As illustrated in Fig. 1, our method differs fundamentally from prior approaches: continuous diffusion models operate in Gaussian space and perform regression-based noise prediction (Fig. 1a), while pseudo-discrete models simulate diffusion through masked-token recovery (Fig. 1b). In contrast, ADD begins from Gaussian-corrupted one-hot vectors and performs true denoising entirely within the one-hot space, using $\arg\max$-based discretization and iterative noise reduction to progressively refine categorical predictions (Fig. 1c).
We evaluate ADD on two tasks—classification and text generation (e.g., image captioning)—as practical benchmarks for discrete generation. Our contribution lies in establishing a discrete-space diffusion process that is both stable and effective, without relying on diffusion in the embedding space or masking-based approximations. The resulting model achieves strong performance with minimal sampling steps and offers a promising direction for generative modeling in categorical and symbolic domains.
Our key contributions are summarized as follows:
  • We propose ADD, an authentic discrete diffusion framework that preserves the essential principles of generative diffusion while operating entirely in the one-hot space.
  • We introduce a timestep-conditioned cross-entropy loss that directly supervises categorical predictions, avoiding the smoothing effects of MSE and ensuring the mutually exclusive nature of discrete outputs.
  • We design an iterative refinement mechanism that starts from Gaussian-corrupted one-hot vectors and progressively denoises them through $\arg\max$-based discretization and reduced-noise feedback.
  • We evaluate ADD on both classification and text generation (e.g., image captioning) as representative discrete generation tasks, achieving excellent performance with minimal sampling steps, and perform comprehensive ablation studies to isolate the contribution of each component.
2 Related Work
Diffusion Models in Continuous Space.
Diffusion models have achieved state-of-the-art performance in continuous generative modeling, particularly in image synthesis. The foundational Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) models data generation as the reverse of a Gaussian noising process, trained with a mean squared error (MSE) loss to predict added noise. Numerous extensions have enhanced efficiency and fidelity: DDIM (Song et al., 2021) accelerates generation via non-Markovian deterministic sampling; ADM (Dhariwal and Nichol, 2021) incorporates classifier guidance and adversarial techniques; and LDM (Rombach et al., 2022) reduces computation through latent-space diffusion. Despite their success, these methods inherently rely on continuous signal representations and quadratic losses, making them poorly suited for discrete data whose outputs lie on a simplex. One line of work attempts to address this by mapping discrete variables into continuous latent spaces via embedding (Gong et al., 2023). However, such approaches often suffer from degraded generation quality and unstable training, as the continuous diffusion process struggles to model sharp, mutually exclusive distributions.
Pseudo-Discrete Diffusion Models.
Extending diffusion models to discrete domains remains an open challenge. A common alternative reframes “discrete diffusion” as a masked token recovery task (Google DeepMind, 2025; Nie et al., 2025; Wu et al., 2025; Ye et al., 2025). In this view, noise is simulated through random masking of tokens, and the model is trained to reconstruct the original input, effectively mirroring masked language modeling as popularized by BERT (Devlin et al., 2019). While such models can be effective in certain language or multimodal applications, they deviate from the formal definition of diffusion (Sohl-Dickstein et al., 2015), lacking a principled forward noising process and stochastic transition dynamics. Our proposed ADD framework addresses this conceptual gap by preserving the defining characteristics of diffusion in a truly discrete setting. Unlike embedding-based or masking-based approaches, ADD operates directly in one-hot space. In the forward process, a clean one-hot vector is perturbed with Gaussian noise under a variance schedule, producing Gaussian-corrupted one-hot vectors. In the reverse process, the model iteratively denoises these vectors: given a corrupted input, it predicts class probabilities, which are discretized via $\arg\max$ into one-hot form. This estimate is supervised by timestep-conditioned cross-entropy alignment with the target one-hot vector, then re-noised with a reduced coefficient for the next iteration. This $\arg\max$-and-re-noise loop constitutes a genuine diffusion mechanism over discrete symbols, supporting both single-token tasks (e.g., classification) and multi-token generation (e.g., text).
Bridging Generative and Discriminative Learning.
There is growing interest in applying diffusion models or generative models to discriminative tasks such as classification (Wang and Torr, 2022; Wang et al., 2022), segmentation, and object detection. Existing methods often use diffusion indirectly—for example, as a feature extractor (Zhu et al., 2025), a synthetic data generator (Li et al., 2023b; Nguyen et al., 2023; Ma et al., 2023; Wu et al., 2023), or a source of attention cues for zero-shot reasoning (Yang et al., 2025; Liu et al., 2024; Ni et al., 2023). Another line explores zero-shot generative classifiers, where classification is formulated as finding the class condition that minimizes diffusion loss (Li et al., 2023a). While promising, these approaches treat discriminative tasks as secondary outcomes of generative modeling, often requiring external guidance or downstream fine-tuning. Recent works (Chen et al., 2025) also apply diffusion models to segmentation tasks and achieve remarkable performance, but these approaches still operate in continuous space and rely on MSE loss. In contrast, ADD directly formulates discrete prediction as diffusion in one-hot space. For single-token tasks such as classification and multi-token tasks such as text generation, denoising itself serves as categorical prediction. This design eliminates the need for handcrafted supervision signals or embedding-space generation, realizing discrete generation through authentic diffusion in symbolic space.
Figure 2: Overview of the proposed framework. (a) Single-token generation for classification. Ground-truth labels are converted into one-hot vectors and perturbed with Gaussian noise during training. The diffusion model iteratively denoises these corrupted vectors back to categorical one-hot outputs through the “to one” operation—implemented as softmax with timestep-conditioned cross-entropy supervision in training and as $\arg\max$ with one-hot vectorization during sampling. Image features provide conditioning signals that guide the denoising process. (b) Conditioning module. The feature extractor encodes the input image into tokens, where some learnable class tokens interacts with image tokens through stacked Transformer layers to yield the conditioning representation $c$, which is injected into diffusion blocks. (c) Multi-token generation for text. ADD extends naturally from single-label classification to sequence generation by applying Gaussian corruption and denoising to each token in a sequence of one-hot vectors. The diffusion blocks predict categorical distributions for all tokens in parallel, enabling efficient iterative refinement of entire sequences into coherent text.
3 Methodology
The goal of this work is to develop an Authentic Discrete Diffusion (ADD) framework that retains the essential characteristics of diffusion models while operating directly in one-hot space. Unlike prior pseudo-discrete approaches that rely on masked-token recovery or diffusion in embedding space, ADD defines both forward and reverse processes directly on categorical one-hot vectors. At every step, predictions are projected back to the one-hot domain, ensuring categorical consistency and enabling stable discrete-space diffusion.
The framework integrates three coordinated components: (i) a forward process that perturbs one-hot labels with Gaussian noise, (ii) a reverse process that iteratively refines predictions during inference, and (iii) a timestep-conditioned cross-entropy objective that supervises training. Conditioning signals (e.g., image features or other modalities) provide semantic guidance but do not replace the discrete generative process. An overview of the pipeline is shown in Fig. 2.
3.1 Background: Diffusion Models
Diffusion models corrupt data progressively with Gaussian noise in a forward process and learn to recover the clean signal through a reverse process. Formally, given data $\mathbf{x}_{0}$, the forward distribution is
$q(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\mathcal{N}\!\left(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}\right),$
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ denotes the variance schedule. The reverse process is parameterized as
$p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})=\mathcal{N}\!\big(\mathbf{x}_{t-1};\,\mu_{\theta}(\mathbf{x}_{t},t),\sigma_{t}^{2}\mathbf{I}\big).$
Standard diffusion models are trained by regressing Gaussian noise with an MSE objective, which is well-suited to continuous domains but fails to respect the mutually exclusive structure of categorical labels.
3.2 Authentic Discrete Diffusion
ADD reformulates diffusion over categorical one-hot vectors. Let $\mathbf{y}_{0}\in\{0,1\}^{K}$ denote a one-hot label over $K$ classes, with $\sum_{k=1}^{K}y_{0}^{(k)}=1$. We also experimented with $\lambda$-hot labels, where $\lambda\in(0,1)$, to make the representations more easily perturbed by Gaussian noise. While this variant led to lower training loss, it did not yield noticeable improvements in accuracy. For this reason, we omit detailed results here and leave further exploration of $\lambda$-hot labels to future work.
Forward process.
At each step $t\in\{1,\dots,T\}$, the one-hot label is perturbed by Gaussian noise:
$q(\mathbf{y}_{t}\mid\mathbf{y}_{0})=\mathcal{N}\!\big(\mathbf{y}_{t};\,\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{0},\,(1-\bar{\alpha}_{t})\mathbf{I}\big),$
producing Gaussian-corrupted vectors $\mathbf{y}_{t}$ that remain close to the categorical simplex.
Reverse process (training).
The denoising network parameterizes a categorical distribution:
$p_{\theta}(\mathbf{y}_{0}\mid\mathbf{y}_{t},c)=\text{Softmax}\!\big(f_{\theta}(\mathbf{y}_{t},t,c)\big),$
where $c$ denotes conditioning signals. Training minimizes the timestep-conditioned cross-entropy loss:
$\mathcal{L}_{\text{CE}}=-\mathbb{E}_{t\sim\mathcal{U}[1,T]}\bar{\alpha}_{t}\sum_{k=1}^{K}y_{0}^{(k)}\log p_{\theta}(y_{0}^{(k)}\mid\mathbf{y}_{t},c).$
The decay coefficient $\bar{\alpha}_{t}$ helps prevent the model from overly relying on the conditions of the conditional diffusion model, thereby promoting the effective learning of the diffusion network.
Reverse process (inference).
At inference, predictions are discretized to remain in the one-hot space:
$\hat{\mathbf{y}}_{0}=\text{onehot}\!\Big(\arg\max_{k}p_{\theta}(y_{0}^{(k)}\mid\mathbf{y}_{t},c)\Big).$
The discretized prediction is re-noised with reduced variance:
$\mathbf{y}_{t-1}\sim\mathcal{N}\!\left(\sqrt{\alpha_{t-1}}\,\hat{\mathbf{y}}_{0},\,(1-\alpha_{t-1})\mathbf{I}\right).$
Iterating this “argmax-and-re-noise” loop progressively sharpens predictions until a clean one-hot label is recovered and we visualize this phenomenon in Fig 6.
3.3 Conditioning with Feature Extractors
ADD incorporates conditional signals for guidance while preserving the discrete diffusion pipeline. For image-conditioned tasks, an encoder $\mathcal{E}$ maps an input $\mathbf{x}$ to tokens:
$\mathbf{z}=\mathcal{E}(\mathbf{x})\in\mathbb{R}^{L\times d}.$
A learnable class token $\mathbf{z}_{\text{cls}}$ is prepended, yielding
$\tilde{\mathbf{z}}=[\mathbf{z}_{\text{cls}};\mathbf{z}_{1},\dots,\mathbf{z}_{L}].$
After $M$ Transformer layers, either the updated class token or the average of other tokens serves as the conditioning vector:
$c=\mathbf{z}_{\text{cls}}^{(M)}\quad\text{or}\quad c=\text{mean}(\mathbf{z}_{\text{1}}^{(M)},\dots,\mathbf{z}_{\text{L}}^{(M)}).$
This vector $c$ is injected into the denoising network, as defined in Section 3.4. To reduce computation and prevent the learned condition from overshadowing the learning of the subsequent diffusion network, we adopt a feature reuse strategy inspired by (Li et al., 2024): (1) Expand each feature batch $K$ times, (2) Sample $K$ distinct timesteps per instance to generate noisy labels, and (3) Inject the same features into all noisy labels. This $K$-fold strategy (with $K=4$ in practice) enables multi-timestep optimization without repeated feature extraction.
3.4 Single and Multiple Token Generation
Classification.
For single-label classification, ADD diffuses and denoises one-hot vectors corresponding to class labels. Conditioning (e.g., image features) guides the denoising process, while timestep-conditioned cross-entropy supervision enforces categorical fidelity. We adopt a simple Transformer to inject conditioning into the diffusion block (see Fig. 2). This setup demonstrates the stability and efficiency of ADD in the simplest categorical setting.
Text generation.
ADD extends naturally to sequences. A sentence is represented as $\mathbf{Y}_{0}=[\mathbf{y}_{0,1},\dots,\mathbf{y}_{0,N}]$, where each $\mathbf{y}_{0,i}$ is a one-hot vector over the vocabulary. The forward process perturbs each token independently:
$q(\mathbf{y}_{t,i}\mid\mathbf{y}_{0,i})=\mathcal{N}\!\big(\mathbf{y}_{t,i};\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{0,i},(1-\bar{\alpha}_{t})\mathbf{I}\big).$
The reverse process predicts distributions in parallel:
$p_{\theta}(\mathbf{y}_{0,i}\mid\mathbf{y}_{t,i},c)=\text{Softmax}\!\big(f_{\theta}(\mathbf{y}_{t,i},t,c)\big).$
Training minimizes the sequence-level loss:
$\mathcal{L}_{\text{text}}=-\sum_{i=1}^{N}\sum_{k=1}^{K}y_{0,i}^{(k)}\log p_{\theta}(y_{0,i}^{(k)}\mid\mathbf{y}_{t,i},c).$
At inference, each token is discretized by $\arg\max$, and iterative refinement sharpens the entire sequence into a coherent sentence. Unlike autoregressive models, ADD denoises all tokens simultaneously, enabling parallel and efficient text generation. Conditioning is injected via self-attention in the diffusion block (see Fig. 2).
4 Experiments
4.1 Experimental Setup
Datasets.
We evaluate ADD on two representative tasks: large-scale image classification and image-conditioned text generation. For classification, we use the ImageNet benchmark (Deng et al., 2009), following the standard data splitting protocol. Specifically, we adopt the $224\times 224$ resolution variant, which contains 1.28 million training images and 50,000 validation images across 1,000 categories. All images are normalized using channel-wise mean values of $(0.485,0.456,0.406)$ and standard deviation values of $(0.229,0.224,0.225)$. For text generation, we conduct experiments on the COCO Captions dataset (Lin et al., 2014), consisting of 82,783 training and 40,504 validation images paired with five human-annotated captions each. Images are preprocessed with standard resizing and cropping, and captions are tokenized into sequences of one-hot vectors over a vocabulary of 10,112 words.
Model Configuration.
Our full architecture contains 111 million trainable parameters. The Transformer encoder dominates the parameter budget (87M) and specializes in hierarchical feature extraction, while the diffusion module contains 24M parameters and is dedicated to discrete-space denoising. This asymmetry reflects our design principle of prioritizing feature quality while maintaining efficient generative refinement. The patchifier resolution is fixed at $224\times 224$ for image classification task and fixed at $256\times 256$ for image captioning task. Detailed layer specifications are provided in the supplementary material.
Training Protocol.
ADD is trained from scratch for 500 epochs using AdamW (Loshchilov and Hutter, 2019). We adopt a base learning rate of $1\!\times\!10^{-4}$ with a weight decay of $0.3$ and effective batch size of $4096$, distributed across $8\times$ NVIDIA A100 80GB GPUs. A warmup of $20$ epochs is followed by cosine learning rate decay (Goyal et al., 2017). Gradient clipping is applied with a global norm of $3.0$, and PyTorch AMP is used for mixed precision training. For COCO captioning, we employed smaller batch sizes and weight decay values.
Infrastructure.
[…]
Evaluation Metrics.
For ImageNet classification, we report Top-1 accuracy on the validation set. For COCO captioning, our primary metric is CLIP Scores computed between image and text features; higher is better. Additionally, we also include qualitative examples to assess fluency and image–text alignment (Fig. 7).
Architecture Design.
ADD integrates a Transformer feature extractor with a diffusion-based categorical denoiser, creating a hybrid architecture optimized for representation learning and generative refinement. Crucially, the same diffusion backbone is shared across tasks: for classification, it denoises single one-hot vectors, while for captioning, it denoises token sequences in parallel.
Our multi-timestep feature reuse strategy further enhances training efficiency. For each encoder-derived embedding, we sample $K=4$ distinct diffusion timesteps per training batch, expanding supervision across multiple noise levels without recomputing features. This yields substantial savings in compute and improves convergence speed.
4.2 Single-Token Generation Task (Classification)
Table 1: Comparison with state-of-the-art methods on ImageNet. All models are trained in an end-to-end manner. Backbone architectures include ViT-Base/16, ViT-Large/16, and ViT-Huge/14 (Dosovitskiy et al., 2020). Results are reported at an input resolution of $224\times 224$, with ViT-H additionally evaluated at $448\times 448$. “*” denotes results reported by (Li et al., 2024). The best result is highlighted in bold.
Methods Resolution Patch #params Top-1 (%)
ViT-Base (standard classification) * $\times224$ $\times16$ 87M 82.3
ViT-Large (standard classification) * $\times224$ $\times16$ 305M 82.6
ViT-Huge (standard classification) * $\times448$ $\times14$ 632M 83.1
ViT-Base (ADD with timestep-conditioned coefficients) $\times224$ $\times16$  [^1] M 82.8
ViT-Base (ADD without timestep-conditioned coefficients) $\times224$ $\times16$  [^1] M 83.0
4.2.1 Comparison with State-of-the-Art Methods
We begin by comparing our approach with state-of-the-art image classification models on ImageNet. Results are summarized in Tab. 1, from which two key observations emerge that highlight the advantages of our method.
ADD vs. Standard Classifiers.
First, we consider a direct comparison against standard classification models. To ensure fairness, we construct a counterpart model by replacing the authentic discrete diffusion module with a conventional classifier head—specifically, a linear layer that outputs classification logits—while keeping all other components identical. Concretely, this corresponds to attaching a linear classification head to the feature extractor shown in Fig. 2. The results in Table 1 are striking: ADD consistently outperforms the standard classifier baseline by a clear margin on the classification task. Our “ViT-Base (ADD)” model achieves a top-1 accuracy of 82.8%, substantially higher than its fair counterpart “ViT-Base (standard classification).” Even more remarkably, “ViT-Base (ADD)” surpasses “ViT-Large (standard classification)” variants that employ nearly three times more parameters (305M vs.  [^1] M). These comparisons strongly validate the effectiveness and efficiency of ADD.
To compare with standard classifiers methods, we implemented ViT-Base and successfully achieved an accuracy proposed in (Li et al., 2024) of 82.3%. We found that ViT-Base reached its highest accuracy around 250 epochs, after which it began to overfit and saw no further improvement. Our ADD method, however, reached the same accuracy as ViT-Base after 300 epochs and showed a trend of continued improvement. Therefore, we increased the number of ADD epochs until convergence. We finally achieved optimal performance of ADD after 500 epoch of training, 82.8%. See Fig. 3 for a detailed comparison.
Figure 3: Comparison of the accuracy of ADD and Classifier in different epochs.We trained ADD and our own state-of-the-art method on ImageNet for 300 epochs and measured the accuracy at each epoch. We found that the state-of-the-art method began to overfit after 250 epochs, while our method showed potential for further improvement even after 300 epochs. This explains why we extend the number of training epochs and ultimately achieved optimal performance (Top-1 score = 82.8) after 500 epochs.
We also tested the classification accuracy of ADD on ImageNet at different sampling steps. We found that ADD can achieve good classification results in about 10 sampling steps, and the best result is achieved after 20 iterations. Details shown as Fig. 5.
Figure 4: Performance Comparison Between MSE Loss and Cross-Entropy Loss. We compared the outputs generated by models trained with the two loss functions. Models trained with cross-entropy loss achieved high accuracy, while those trained with MSE loss showed unsatisfactory performance.
Figure 5: Performance test of ADD at different sampling steps. We conducted experiments with and without timestep-conditioned coefficients respectively, and tested the classification accuracy of ADD on ImageNet at different sampling steps. We found that ADD can achieve good classification results in about 10 sampling steps, and the best result is achieved after 20 iterations. However, without the timestep-conditioned coefficients, the iterative process exhibits slightly higher accuracy
4.2.2 Ablation Analysis
We conduct rigorous ablations to examine the design choices of our framework. All variants share the same Transformer backbone (ViT-Base(Dosovitskiy et al., 2021),  [^1] M parameters) and training configuration. All controlled experiments were conducted under the condition of training for a full 500 epochs, where the model reached a converged state. The results demonstrate that by integrating several techniques, we achieve the performance presented in Table 1.
Table 2: Ablation study. We tested the accuracy of ADD using various training and generation strategies. We found that using timestep-conditioned cross entropy loss, classifier-free guidance, and “arg max + one hot” configuration achieved the best performance for ADD.
Methods Epoch Top-1 (%)
ADD (discrete diffusion with regressive loss) 400 0.13
ADD (discrete diffusion with timestep-conditioned cross entropy loss) 400 82.72
ADD (without classifier-free guidance) 500 82.36
ADD (with classifier-free guidance) 500 82.82
ADD (with timestep-conditioned coefficients) 500 82.82
ADD (without timestep-conditioned coefficients) 500 82.96
ADD (sampling with softmax) 500 82.35
ADD (sampling with $\arg\max$ + one-hot”) 500 82.82
MSE Loss vs. Cross-Entropy Alignment Loss.
We first assess the necessity of timestep-conditioned cross-entropy alignment loss. In training diffusion models, regressive losses (e.g., MSE, $\mathcal{L}_{1}$) are typically used. In contrast, ADD employs timestep-conditioned cross-entropy to enforce prediction of $x_{0}$ (also denoted $x_{\text{start}}$). To test its importance, we replace timestep-conditioned cross-entropy with a regressive loss, training the model to predict noise as in conventional diffusion. Table 2 and Fig. 4 show that this substitution causes catastrophic degradation (82.73% $\rightarrow$ 0.13%), confirming that timestep-conditioned cross-entropy is indispensable for effective discrete-space diffusion. A comparison of the iterative processes for the two loss functions can be found in Fig. 4.
The Timestep-Conditioned Coefficient in the Cross-Entropy Loss
The effectiveness of the timestep-conditioned coefficients in enhancing the iterative process is experimentally demonstrated in Fig. 5. With the introduction of the timestep-dependent coefficient, the iterative process during generation significantly enhances the model’s ability to progressively refine its predictions, achieving remarkably high classification accuracy after sampling 10 steps. Meanwhile, it shows that removing this coefficient leads to higher accuracy but more modest improvements throughout the iterative process.
Classifier-Free Guidance.
Classifier-free guidance has proven crucial in continuous diffusion models for conditional generation, as it strengthens alignment between generated outputs and conditioning signals. To verify whether this property holds in our discrete setting, we remove classifier-free guidance from ADD while keeping all other components fixed. The performance drops notably (82.82% $\rightarrow$ 82.36%), as shown in Table 2. This confirms that classifier-free guidance plays an equally indispensable role in discrete diffusion, ensuring semantic consistency with conditioning inputs. Importantly, this result demonstrates that ADD retains the core mechanics of standard diffusion frameworks, thereby reinforcing our claim that ADD is an authentic discrete diffusion model.
Sampling Strategy.
Finally, we compare two “to-one” operations used during sampling: (i) softmax-based sampling and (ii) $\arg\max$ followed by one-hot projection. As illustrated in Table 2, both strategies perform competitively, though $\arg\max$ + one-hot projection yields a slight but consistent advantage. Consequently, unless otherwise specified, all sampling results in this paper use $\arg\max$ + one-hot projection as the default strategy. For the specific accuracy changes resulting from different iteration settings during sampling, refer to Fig. 5. Futhermore, Fig. 6 demonstrates that the output distribution of our model gradually approaches a one-hot distribution as the iterations proceed.
Figure 6: The distribution change of model output with sampling steps. To verify that the model can gradually generate one-hot labels through denoising, we visualized the output of the model at each sampling step. The results show that when the sampling step is set to 20, the model is able to gradually denoise pure Gaussian noise into an approximate one-hot vector.
4.3 Multi-Token Generation Task (Text Generation)
Figure 7: Quality comparison of ADD and PDD. The ground-truth captions are shown alongside captions generated by ADD and PDD for representative COCO examples. ADD produces fluent and semantically accurate captions that align well with both the image content and ground-truth descriptions, whereas PDD often generates ungrammatical or semantically inconsistent text. This highlights the superiority of authentic discrete diffusion (ADD) over pseudo discrete diffusion (PDD) in multi-token text generation.
Table 3: CLIP Scores metric testing and comparison. CLIP Scores measure the cosine similarity between the normalized features of image-text pairs. In this part, we validated the similarity for three distinct types of such pairs.
Selection of Text for Computation CLIP Scores.
Ground-Truth captions 0.30
Shuffled captions 0.16
PDD-Generated captions 0.18
ADD-Generated captions 0.25
We further investigate the potential of ADD in multi-token generation tasks, with a particular focus on image captioning as a representative text generation benchmark. Unlike single-label classification, captioning requires the model to generate coherent sequences of tokens conditioned on visual inputs, thereby testing its ability to model discrete sequential dependencies. This task is especially challenging for diffusion-based approaches, as it requires capturing both the syntactic structure of natural language and the semantic alignment with images.
A major difficulty in this setting is the relatively limited scale of available text supervision in datasets such as COCO, which provides only short captions per image. In contrast, state-of-the-art large language models are typically trained on massive corpora with billions of tokens using extensive computational resources. Consequently, training a competitive captioning model directly on COCO represents a stringent test of efficiency and adaptability. To address the scarcity of text tokens, we utilized CLIP Scores as the evaluation metric. This metric employs CLIP to extract image and text features, and calculates the similarity between these two features, providing an well-suited evaluation of semantic relevance. In addition, we present qualitative examples to illustrate the fluency and semantic accuracy of generated captions.
For comparison, we benchmark our ADD against the masked diffusion framework, which we refer to as the “pseudo” discrete diffusion (PDD) model. PDD generates tokens by predicting masked entries in a partially observed sequence, but unlike ADD, it does not operate in a fully discrete diffusion process and therefore lacks consistent alignment with categorical token spaces.
To evaluate the semantic relevance between the captions and the corresponding images quantitatively, we introduced a pre-trained CLIP model in the experimental section to extract features from both text and images and compute their similarity in Table 3. We evaluated the CLIP Scores for several combinations: genuine image-text pairs, mismatched pairs (where the text is grammatically correct but largely unrelated to the image), PDD-generated text paired with images, and ADD-generated text paired with images. The results clearly demonstrate that the captions predicted by ADD exhibit strong semantic alignment with the images.
Beyond numerical evaluation, qualitative comparisons are provided in Fig. 7. Captions generated by ADD exhibit grammatical correctness, semantic coherence, and close alignment with both the input image content and human-annotated ground-truth captions. In contrast, outputs from PDD frequently suffer from broken syntax and poor semantic fidelity, often producing phrases that are disconnected from the visual scene. These observations reinforce our claim that ADD, by operating directly in the one-hot label space with timestep-conditioned cross-entropy–based training, preserves both the discrete structure of text tokens and their semantic alignment with conditioning inputs.
Taken together, these results highlight two key insights. First, authentic discrete diffusion substantially improves text generation quality over pseudo-discrete alternatives. Second, our findings suggest that ADD can serve as a foundation for broader applications in multi-token generation, bridging the gap between discrete modeling in vision tasks and natural language generation. This dual capability further strengthens the generality of ADD as a unified framework for discrete generative modeling.
5 Conclusion
We presented Authentic Discrete Diffusion (ADD), a new framework that extends diffusion modeling to categorical domains while preserving its core generative principles. Unlike prior pseudo-discrete approaches that rely on embeddings or masking-based approximations, ADD operates directly in the one-hot space through a combination of cross-entropy alignment loss, $\arg\max$-based discretization, and iterative denoising. Our experiments on classification and text generation demonstrate that ADD achieves strong performance with minimal sampling steps, significantly outperforming pseudo-discrete baselines in both quantitative metrics (e.g., accuracy and perplexity) and qualitative evaluations. Ablation studies further highlight the indispensability of each component, confirming that cross-entropy alignment loss and classifier-free guidance are critical to making discrete diffusion both stable and effective. Overall, ADD establishes a principled and efficient pathway for applying diffusion models to symbolic data. We believe this work provides a foundation for future research on discrete generative modeling, with potential extensions to large-scale language tasks, multimodal reasoning, and structured prediction problems.
6 Future Work
While Authentic Discrete Diffusion (ADD) establishes a principled framework for discrete-space generative modeling, several promising directions remain open for future exploration.
Theoretical Foundations.
A natural extension of this work is to formalize the theoretical underpinnings of ADD. Current results demonstrate empirical stability and effectiveness, yet deeper analysis is needed to rigorously characterize its convergence behavior, sample complexity, and generalization bounds. Establishing guarantees on the approximation quality of $\arg\max$-based denoising, as well as the statistical efficiency of noise-schedule–aware cross-entropy weighting, would provide a stronger foundation for understanding ADD as a discrete analog of continuous diffusion. Moreover, connecting ADD to information-theoretic principles—such as entropy reduction in categorical spaces—could yield general laws governing discrete diffusion processes.
Scaling Laws.
Future applications of ADD span language, multimodal, and embodied domains. In large language models, ADD offers a discrete diffusion alternative to autoregressive training, with the potential to mitigate exposure bias, improve long-sequence stability, and enable efficient multi-token prediction. Extending ADD to multimodal LLMs could unify discrete-space denoising across text, vision, and structured categorical data, fostering categorical consistency and interpretability in cross-modal reasoning tasks such as grounding and retrieval. Beyond language and vision, ADD also provides a pathway toward large physical world models, enabling scalable and interpretable systems that integrate reasoning, perception, and action for embodied AI.